OpenStudy (anonymous):

got a couple questions Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14 then Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3 Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x

OpenStudy (anonymous):

answer for last question y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (aum):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). Same method as we did last time. If you learn it you can answer all these on your own. 3x - y = -3 Put this in the form y = mx + b Add y to both sides: 3x = y - 3 add 3 to both sides: 3x + 3 = y or y = 3x + 3 compare this to y = mx + b where m is the slope. m = 3. Parallel lines have the same slope. So the slope of the line we are looking for is 3. y = 3x + b It passes through the point (4,-2) put x = 4 and y = -2 and solve for b -2 = 3(4) + b -2 = 12 + b b = -14 y = 3x - 14

OpenStudy (anonymous):

ok can u stay just to see if i do the next one right

OpenStudy (anonymous):

i got d bc i add both sides like u said got 4.5 then so i got the x then for b i got 3

OpenStudy (anonymous):

did i do it right